OpenStudy (anonymous):

what is 32 over 10 as a mixed number in simplest form?

jimthompson5910 (jim_thompson5910):

what is 32/10 in decimal form?

OpenStudy (anonymous):

3.2

jimthompson5910 (jim_thompson5910):

now multiply the decimal portion 0.2 by 10 to get ______

OpenStudy (anonymous):

0.02

jimthompson5910 (jim_thompson5910):

you went the wrong way

jimthompson5910 (jim_thompson5910):

it should be 0.2 * 10 = 2

OpenStudy (anonymous):

use the division algorithm to rewrite 32 => 32 = 3*10 + 2 so as a mixed number \[\frac{ 32 }{ 10 }=\frac{ 3\cdot 10 }{10 }+\frac{ 2 }{10 }= 3+\frac{ 1 }{5 }=3\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

Thank you you'r the best @pgpilot326

OpenStudy (anonymous):

*your

jimthompson5910 (jim_thompson5910):

32/10 = 3 remainder 2 so that's why \[\large \frac{32}{10} = 3\frac{2}{10}\] that turns into \[\large \frac{32}{10} = 3\frac{1}{5}\]

OpenStudy (anonymous):

OKay thank to all of you

OpenStudy (anonymous):

I have never learned it that way.... 32 over 10 each can be divided by 2 16 over 5 5 goes into 16 three times with a remainder of 1 3 and 1/5

OpenStudy (anonymous):

you're welcome!

OpenStudy (debbieg):

Jim's approach works, but only because you have a terminating decimal. The long division approach is more "universal". Dividend/divisor=quotient + remainder/divisor